# SKD-TSTSAN: Three-Stream Temporal-Shift Attention Network Based on Self-Knowledge Distillation for Micro-Expression Recognition

### v1

@@ Abstract
Micro-expressions (MEs) are subtle facial movements that occur spontaneously when people try to conceal the real emotions. Micro-expression recognition (MER) is crucial in many fields, including criminal analysis and psychotherapy. However, MER is challenging since MEs have low intensity and ME datasets are small in size. To this end, a three-stream temporal-shift attention network based on self-knowledge distillation (SKD-TSTSAN) is proposed in this paper. Firstly, to address the low intensity of ME muscle movements, we utilize learning-based motion magnification modules to enhance the intensity of ME muscle movements. Secondly, we employ efficient channel attention (ECA) modules in the local-spatial stream to make the network focus on facial regions that are highly relevant to MEs. In addition, temporal shift modules (TSMs) are used in the dynamic-temporal stream, which enables temporal modeling with no additional parameters by mixing ME motion information from two different temporal domains. Furthermore, we introduce self-knowledge distillation (SKD) into the MER task by introducing auxiliary classifiers and using the deepest section of the network for supervision, encouraging all blocks to fully explore the features of the training set. Finally, extensive experiments are conducted on four ME datasets: CASME II, SAMM, MMEW, and CAS(ME)3. The experimental results demonstrate that our SKD-TSTSAN outperforms other existing methods and achieves new state-of-the-art performance. Our code will be available at https://github.com/GuanghaoZhu663/SKD-TSTSAN.
@@ 1 Introduction
Facial expression (FE) is a common means of nonverbal communication and is crucial for understanding human mental states and intentions [1]. Facial expressions can generally be categorized into two classes: macro-expressions and micro-expressions (MEs). The major variations between them are intensity and duration. Macro-expressions typically have a duration of 0.5-4 seconds and involve muscle movements covering a large area of the face, which can easily be recognized by humans [2]. In contrast, MEs reflect hidden emotions that come out spontaneously when people strive to hide their true emotions. It usually lasts only 1/25 to 1/3 of a second with muscle movements in a minor region of the face [3], which is difficult to recognize, and only an expert with extensive training can distinguish MEs. MEs are significant in many fields, such as crime analysis [4], emotional communication, and clinical diagnosis. Therefore, researchers are striving to utilize computer vision techniques for automatic micro-expression recognition (MER).
@@
Early MER methods are based on manual feature extraction, such as Local Binary Pattern histograms from Three Orthogonal Planes (LBP-TOP) [5] and Main Directional Mean Optical-flow (MDMO) [6]. These methods largely rely on manually designed extractors and complicated parameter adjustments. The extracted features are mainly low-level features [1], therefore it is difficult to recognize the subtle movements in MEs with high accuracy. With the proposal of a series of ME datasets [7, 8, 9, 10] and the development of deep learning, researchers began to use CNN to extract spatial and temporal features of MEs. For example, Song et al. [11] proposed a three-stream convolutional neural network (TSCNN) to learn the spatiotemporal information of MEs. Xia et al. [12] trained two dual-stream baseline models called MiNet and MaNet using the ME dataset and macro-expression dataset. By introducing two auxiliary tasks between the two models, MiNet can effectively extract both the dynamic patterns of the muscle motion and the static texture of the face. Although these methods can extract ME features and achieve promising results, there are still two major challenges in MER: the low intensity of MEs and overfitting on small ME datasets [9]. Since MEs have low intensity, networks need to have strong representation ability, which generally means more convolutional layers and computational costs. However, complex models usually overfit on small-scale ME datasets. Therefore, in order to mine subtle ME motions from small ME datasets, more effective blocks need be introduced.
@@
To address the above challenges, we propose a three-stream temporal-shift attention network based on self-knowledge distillation called SKD-TSTSAN in this paper. Firstly, we use the TSCNN [11] as our baseline, which consists of three streams, namely static-spatial stream (S-stream), local-spatial stream (L-stream), and dynamic-temporal stream (T-stream). The inputs of the three streams are the whole facial image, the concatenation of $n\times n$ localized facial images, and optical flow images, respectively. The three recognition streams are finally fused into a fully connected layer for MER tasks. Targeting the low intensity of MEs, we add pre-trained magnification networks [13] to the head of each convolutional stream based on transfer learning. For an end-to-end framework, a magnification network without an encoder is employed. Since ME only appears in local regions of the face, we incorporate efficient channel attention (ECA) modules [14] into the L-stream to encourage the network to focus on significant local regions.
@@
Compared with 2D CNNs, 3D CNNs can mine the spatiotemporal information of videos more effectively [1]. However, 3D CNNs generally have a large number of parameters and are easily overfitted on small-scale ME datasets [1]. To capture temporal relations in ME sequences and maintain the complexity of 2D CNNs, we introduce temporal shift modules (TSMs) [15] in T-stream. In TSM, part of the optical flow features’ channels are shifted along the temporal dimension to realize the information exchange. Moreover, to further enhance the performance of the lightweight network, the macro-expression dataset CK+ [16] is employed for pre-training, and self-knowledge distillation (SKD) [17] is incorporated. Additional convolutional and fully connected layers are added to form multiple auxiliary classifiers (ACs). By using labels and the model’s final output to train the ACs, we enable all of the model’s blocks to fully explore the training set’s features.
@@ 2 Related Work
In this section, we initially introduce current micro-expression recognition methods, including hand-crafted methods and deep learning methods. Then, since our proposed SKD-TSTSAN uses temporal shift modules (TSMs) [15] for temporal modeling in MER and incorporates the self-knowledge distillation (SKD) method [17], we also briefly introduce temporal modeling and knowledge distillation.
@@ 2.1.1 Hand-crafted methods
Hand-crafted methods commonly extract manually designed appearance-based ME features and then provide them to classic classifiers for emotion recognition, including the K nearest neighbor (KNN) classifier [18, 19] and support vector machine (SVM) [20, 21]. Pfister et al. [5] employed a temporal interpolation model (TIM) to obtain a sufficient amount of frames, and then utilized the LBP-TOP as the texture descriptor for classification. Since then, many LBP-TOP-based enhanced MER methods have been proposed. For example, Wang et al. [22] proposed Local Binary Patterns with Six Intersection Points (LBP-SIP) descriptor, which replaces the three orthogonal planes containing redundant points with intersecting lines of them. Only six different neighboring points around the center point are used to obtain the spatiotemporal LBP patterns, which improves the computational efficiency. In addition to LBP-TOP features, optical flow features, which reflect the motion patterns of moving objects, are also significant in ME research [9]. Liu et al. [6] proposed an MDMO feature that only considers mean optical flow in the main direction, thus having fewer feature dimensions. In conclusion, although these approaches offer strong interpretability, they depend on specialized knowledge and extensive parameter adjustments [9]. Besides, the extracted features are relatively shallow, which is not conducive to recognizing subtle movements in ME.
@@ 2.1.2 Deep-learning methods
Deep learning-based MER methods typically combine feature extraction and classification, achieving SOTA predictive performance in recent years. Since optical flow can reflect object motion by detecting changes in pixel intensity between two frames, it has been widely applied in MER tasks. For example, Wang et al. [23] presented a hierarchical transformer network (HTNet), which first calculates the optical flow between the onset frame and the apex frame, and divides the face into four different facial regions. Then, the transformer layers capture fine-grained features within the local regions, and aggregation blocks facilitate interaction between different regions.
Due to the low intensity of ME, some researchers have utilized motion magnification techniques to enhance muscle movements in ME sequences [24, 25, 26], such as Eulerian Motion Magnification (EMM). Traditional motion magnification techniques relying on hand-crafted filters often result in noise and excessive blurring. Therefore, Oh et al. [13] proposed a learning-based motion magnification network that directly learns filters from data, achieving high-quality magnification results. Subsequently, learning-based motion magnification algorithm has been widely used in MER tasks. For example, Wei et al. [26] employed a set of different amplification factors (AFs) to amplify ME movements and used an attention mechanism to assign varying weights to these AFs, adaptively focusing on the appropriate amplification level.
The ME datasets are relatively small in scale, while macro-expression datasets are large and share some common features, such as similar Action Units (AUs) when expressing emotions [9]. Therefore, using transfer learning to transfer knowledge from macro-expressions to MEs can typically enhance the performance of MER methods. For example, Wang et al. [27] employed a two-step transfer learning approach. Firstly, they pre-trained the CNN using a large-scale dataset of general expressions. Then, they fine-tuned the pre-trained network by treating each frame of the ME video clips as a sample. This approach leverages the high relevance between the source and target domains, ensuring improved performance.
@@ 2.2 Temporal Modeling
Using 3D CNNs is a direct approach for temporal modeling. Li et al. [28] proposed a 3D flow-based CNN composed of three data stream subnetworks. 3D convolutional kernels were employed to extract spatiotemporal feature information. However, using deep 3D CNNs significantly increases computational costs. Thus, Xie et al. [29] adopted a lightweight 3D CNN backbone called STPNet to extract spatiotemporal features, where the 3D convolutional filter is decomposed into 2D spatial convolution and 1D temporal convolution. Wu et al. [30] proposed a network for MER that combines 2D CNNs and 3D CNNs. This network uses 3D convolutional kernels with varying settings to extract features in different temporal domains and then utilizes 2D CNN blocks to extract spatial features.
Another approach for modeling temporal relations is using 2D CNN + post-hoc fusion. Khor et al. [31] proposed an Enriched Long-term Recurrent Convolutional Network (ELRCN), where the ME features are fed into a Long Short-Term Memory (LSTM) network [32] to learn the temporal sequence information of MEs. However, 3D CNN-based methods have high computational requirements and tend to overfit on limited ME datasets, while 2D CNN + post-hoc fusion methods cannot capture low-level information that is lost in the feature extraction process. Therefore, we introduce temporal shift modules (TSMs) [15] into the dynamic-temporal stream of our three-stream network to capture the temporal relationships with no additional parameters.
@@ 2.3 Knowledge Distillation
Knowledge distillation is an effective approach for compressing models. The key is to make a compact student model approximate an over-parameterized teacher model, thereby achieving significant performance improvements [33]. Using a shallow student model instead of a teacher model enables compression and acceleration. Sun et al. [34] employed a pre-trained network for AU detection as the teacher network, extracting knowledge from AUs and transferring it to a shallow student network. However, in traditional
@@
two-stage distillation, designing an appropriate teacher model is challenging. Moreover, excellent student models that outperform teacher models are rare due to inefficient knowledge transfer [17].
To address the shortcomings of traditional distillation, we employ the self-distillation framework proposed in [17], which adds multiple auxiliary classifiers (AC) to the model. The deepest part and the shallow part with classifiers are respectively treated as the teacher model and the student model, achieving single-stage self-knowledge distillation (SKD). Experimental results show that using SKD improves network performance while maintaining low complexity.
3 Method
The framework of our SKD-STSTAN is illustrated in Fig. 1. Firstly, the onset and apex frames of ME sequences as well as the motionless and real optical flow images are fed into the motion magnification modules to obtain the amplified shape representation, which is discussed in detail in Section 3.1. Then, the magnified apex frame representation is input into the static-spatial stream (S-stream) to obtain the contour and appearance information of the whole face. The representation of the magnified local blocks is input into the local-spatial stream (L-stream) for the local ME feature extraction. The efficient channel attention (ECA) module [14] used in the L-stream is introduced in Section 3.2. The amplified optical flow maps are input into the dynamic-temporal stream (T-stream), and the temporal shift module (TSM) [15] used in the T-stream is introduced in Section 3.3. Finally, the outputs of three recognition streams are fused into a fully connected layer to get the classification results. Furthermore, we add multiple auxiliary classifiers (ACs) in the network, leveraging SKD to enable all blocks to fully learn the features of the training set, which is discussed in Section 3.4.
@@ 3.1 Motion Magnification Module
Due to the low intensity of ME, we enhance muscle movements based on transfer learning using the motion magnification network proposed by Oh et al. [13] To achieve an end-to-end framework and avoid noise in magnified images reconstructed in the decoder, our motion magnification module does not use a decoder. Given that each stream in our three-stream network has different inputs, we add separate motion magnification modules for each stream. Specifically, the input to the motion magnification network includes two frames with small motion displacement. For an ME sequence, the onset frame is the first frame where the ME begins, and the apex frame is the frame where the facial expression reaches maximum intensity, conveying significant emotional information. Therefore, the onset and apex frames are employed as inputs for the motion magnification module of the S-stream, and the concatenation of their $n\times n$ grid divisions as inputs for the L-stream’s motion magnification module. Particularly, since the T-stream extracts temporal information from optical flow fields, we utilize the optical flow corresponding to no motion and the optical flow between the onset and apex frames (or: between the apex and offset frames) as inputs for the motion magnification module in the T-stream.
Additionally, according to [24], shape representation often reflects boundaries with geometric properties, while texture representation usually represents color properties. Therefore, shape representation is more helpful for the network to learn the relationships of muscle movements, whereas texture representation provides little useful information for ME. Moreover, the manipulator in the motion magnification module works by calculating the difference between two shape representations and multiplying it by an amplification factor, without using the texture representation, as shown below:

### v2

# Three-Stream Temporal-Shift Attention Network Based on Self-Knowledge Distillation for Micro-Expression Recognition

@@ Abstract
Micro-expressions are subtle facial movements that occur spontaneously when people try to conceal real emotions. Micro-expression recognition is crucial in many fields, including criminal analysis and psychotherapy. However, micro-expression recognition is challenging since micro-expressions have low intensity and public datasets are small in size. To this end, a three-stream temporal-shift attention network based on self-knowledge distillation called SKD-TSTSAN is proposed in this paper. Firstly, to address the low intensity of muscle movements, we utilize learning-based motion magnification modules to enhance the intensity of muscle movements. Secondly, we employ efficient channel attention modules in the local-spatial stream to make the network focus on facial regions that are highly relevant to micro-expressions. In addition, temporal shift modules are used in the dynamic-temporal stream, which enables temporal modeling with no additional parameters by mixing motion information from two different temporal domains. Furthermore, we introduce self-knowledge distillation into the micro-expression recognition task by introducing auxiliary classifiers and using the deepest section of the network for supervision, encouraging all blocks to fully explore the features of the training set. Finally, extensive experiments are conducted on four public datasets: CASME II, SAMM, MMEW, and CAS(ME)3. The experimental results demonstrate that our SKD-TSTSAN outperforms other existing methods and achieves new state-of-the-art performance. Our code will be available at https://github.com/GuanghaoZhu663/SKD-TSTSAN.
@@ 1 Introduction
Facial expression (FE) is a common means of nonverbal communication and is crucial for understanding human mental states and intentions (Li et al., 2022c, ). Facial expressions can generally be categorized into two classes: macro-expressions and micro-expressions (MEs). The major variations between them are intensity and duration. Macro-expressions typically have a duration of 0.5-4 seconds and involve muscle movements covering a large area of the face, which can easily be recognized by humans (Matsumoto and Hwang,, 2011). In contrast, MEs reflect hidden emotions that come out spontaneously when people strive to hide their true emotions. It usually lasts only 1/25 to 1/3 of a second with muscle movements in a minor region of the face (Ekman,, 2009), which is difficult to recognize, and only an expert with extensive training can distinguish MEs. MEs are significant in many fields, such as crime analysis (Polikovsky et al.,, 2009), emotional communication, and clinical diagnosis. Therefore, researchers are striving to utilize computer vision techniques for automatic micro-expression recognition (MER).
@@
Early MER methods are based on manual feature extraction, such as Local Binary Pattern histograms from Three Orthogonal Planes (LBP-TOP) (Pfister et al.,, 2011) and Main Directional Mean Optical-flow (MDMO) (Liu et al.,, 2015). These methods largely rely on manually designed extractors and complicated parameter adjustments. The extracted features are mainly low-level features (Li et al., 2022c, ), therefore it is difficult to recognize the subtle movements in MEs with high accuracy. With the proposal of a series of ME datasets (Yan et al.,, 2014; Davison et al.,, 2016; Ben et al.,, 2021; Li et al., 2022b, ) and the development of deep learning, researchers began to use CNN to extract spatial and temporal features of MEs. For example, Song et al., (2019) proposed a three-stream convolutional neural network (TSCNN) to learn the spatiotemporal information of MEs. Xia and Wang, (2021) trained two dual-stream baseline models called MiNet and MaNet using the ME dataset and macro-expression dataset. By introducing two auxiliary tasks between the two models, MiNet can effectively extract both the dynamic patterns of the muscle motion and the static texture of the face. Although these methods can extract ME features and achieve promising results, there are still two major challenges in MER: the low intensity of MEs and overfitting on small ME datasets (Ben et al.,, 2021). Since MEs have low intensity, networks need to have strong representation ability, which generally means more convolutional layers and computational costs. However, complex models usually overfit on small-scale ME datasets. Therefore, in order to mine subtle ME motions from small ME datasets, more effective blocks need be introduced.
@@
To address the above challenges, we propose a three-stream temporal-shift attention network based on self-knowledge distillation called SKD-TSTSAN in this paper. Firstly, we use the TSCNN (Song et al.,, 2019) as our baseline, which consists of three streams, namely static-spatial stream (S-stream), local-spatial stream (L-stream), and dynamic-temporal stream (T-stream). The inputs of the three streams are the whole facial image, the concatenation of $n\times n$ localized facial images, and optical flow images, respectively. The three recognition streams are finally fused into a fully connected layer for MER tasks. Targeting the low intensity of MEs, we add pre-trained magnification networks (Oh et al.,, 2018) to the head of each convolutional stream based on transfer learning. For an end-to-end framework, a magnification network without an encoder is employed. Since ME only appears in local regions of the face, we incorporate efficient channel attention (ECA) modules (Wang et al.,, 2020) into the L-stream to encourage the network to focus on significant local regions.
@@
Compared with 2D CNNs, 3D CNNs can mine the spatiotemporal information of videos more effectively (Li et al., 2022c, ). However, 3D CNNs generally have a large number of parameters and are easily overfitted on small-scale ME datasets (Li et al., 2022c, ). To capture temporal relations in ME sequences and maintain the complexity of 2D CNNs, we introduce temporal shift modules (TSMs) (Lin et al.,, 2019) in T-stream. In TSM, part of the optical flow features’ channels are shifted along the temporal dimension to realize the information exchange. Moreover, to further enhance the performance of the lightweight network, the macro-expression dataset CK+ (Lucey et al.,, 2010) is employed for pre-training, and self-knowledge distillation (SKD) (Zhang et al.,, 2019) is incorporated. Additional convolutional and fully connected layers are added to form multiple auxiliary classifiers (ACs). By using labels and the model’s final output to train the ACs, we enable all of the model’s blocks to fully explore the training set’s features.
@@ 2 Related Work
In this section, we initially introduce current micro-expression recognition methods, including hand-crafted methods and deep learning methods. Then, since our proposed SKD-TSTSAN uses temporal shift modules (TSMs) (Lin et al.,, 2019) for temporal modeling in MER and incorporates the self-knowledge distillation (SKD) method (Zhang et al.,, 2019), we also briefly introduce temporal modeling and knowledge distillation.
@@ 2.1.1 Hand-crafted methods
Hand-crafted methods commonly extract manually designed appearance-based ME features and provide them to classic classifiers for emotion recognition, including the K nearest neighbor (KNN) classifier (Chaudhry et al.,, 2009; Zhang et al.,, 2017) and support vector machine (SVM) (Wang et al., 2015a, ; Xu et al.,, 2017). Pfister et al., (2011) employed a temporal interpolation model (TIM) to obtain a sufficient amount of frames, and then utilized the LBP-TOP as the texture descriptor for classification. Since then, many LBP-TOP-based enhanced MER methods have been proposed. For example, Wang et al., 2015b proposed Local Binary Patterns with Six Intersection Points (LBP-SIP) descriptor, which replaces the three orthogonal planes containing redundant points with intersecting lines of them. Only six different neighboring points around the center point are used to obtain the spatiotemporal LBP patterns, which improves the computational efficiency. In addition to LBP-TOP features, optical flow features, which reflect the motion patterns of moving objects, are also significant in ME research (Ben et al.,, 2021). Liu et al., (2015) proposed an MDMO feature that only considers mean optical flow in the main direction, thus having fewer feature dimensions. In conclusion, although these approaches offer strong interpretability, they depend on specialized knowledge and extensive parameter adjustments (Ben et al.,, 2021). Besides, the extracted features are relatively shallow, which is not conducive to recognizing subtle movements in ME.
@@ 2.1.2 Deep-learning methods
Deep learning-based MER methods typically combine feature extraction and classification, achieving SOTA predictive performance in recent years. Since optical flow can reflect object motion by detecting changes in pixel intensity between two frames, it has been widely applied in MER tasks. For example, Wang et al., (2023) presented a hierarchical transformer network (HTNet), which first calculates the optical flow between the onset frame and the apex frame, and divides the face into four different facial regions. Then, the transformer layers capture fine-grained features within the local regions, and aggregation blocks facilitate interaction between different regions.
Due to the low intensity of ME, some researchers have utilized motion magnification techniques to enhance muscle movements in ME sequences (Lei et al.,, 2020; Zhang et al.,, 2022; Wei et al.,, 2022), such as Eulerian Motion Magnification (EMM). Traditional motion magnification techniques relying on hand-crafted filters often result in noise and excessive blurring. Therefore, Oh et al., (2018) proposed a learning-based motion magnification network that directly learns filters from data, achieving high-quality magnification results. Subsequently, learning-based motion magnification algorithm has been widely used in MER tasks. For example, Wei et al., (2022) employed a set of different amplification factors (AFs) to amplify ME movements and used an attention mechanism to assign varying weights to these AFs, adaptively focusing on the appropriate amplification level.
The ME datasets are relatively small in scale, while macro-expression datasets are large and share some common features, such as similar Action Units (AUs) when expressing emotions (Ben et al.,, 2021). Therefore, using transfer learning to transfer knowledge from macro-expressions to MEs can typically enhance the performance of MER methods. For example, Wang et al., (2018) employed a two-step transfer learning approach. Firstly, they pre-trained the CNN using a large-scale dataset of general expressions. Then, they fine-tuned the pre-trained network by treating each frame of the ME video clips as a sample. This approach leverages the high relevance between the source and target domains, ensuring improved performance.
@@ 2.2 Temporal Modeling
Using 3D CNNs is a direct approach for temporal modeling. Li et al., (2019) proposed a 3D flow-based CNN composed of three data stream subnetworks. 3D convolutional kernels were employed to extract spatiotemporal feature information. However, using deep 3D CNNs significantly increases computational costs. Thus, Xie et al., (2020) adopted a lightweight 3D CNN backbone called STPNet to extract spatiotemporal features, where the 3D convolutional filter is decomposed into 2D spatial convolution and 1D temporal convolution. Wu and Guo, (2021) proposed a network for MER that combines 2D CNNs and 3D CNNs. This network uses 3D convolutional kernels with varying settings to extract features in different temporal domains and then utilizes 2D CNN blocks to extract spatial features.
Another approach for modeling temporal relations is using 2D CNN + post-hoc fusion. Khor et al., (2018) proposed an Enriched Long-term Recurrent Convolutional Network (ELRCN), where the ME features are fed into a Long Short-Term Memory (LSTM) network (Gers et al.,, 2000) to learn the temporal sequence information of MEs. However, 3D CNN-based methods have high computational requirements and tend to overfit on limited ME datasets, while 2D CNN + post-hoc fusion methods cannot capture low-level information that is lost in the feature extraction process. Therefore, we introduce temporal shift modules (TSMs) (Lin et al.,, 2019) into the dynamic-temporal stream of our three-stream network to capture the temporal relationships with no additional parameters.
@@ 2.3 Knowledge Distillation
Knowledge distillation is an effective approach for compressing models. The key is to make a compact student model approximate an over-parameterized teacher model, thereby achieving significant performance improvements (Hinton et al.,, 2015). Using a shallow student model instead of a teacher model enables compression and acceleration. Sun et al., (2020) employed a pre-trained network for AU detection as the teacher network, extracting knowledge from AUs and transferring it to a shallow student network. However, in traditional two-stage distillation, designing an appropriate teacher model is challenging. Moreover, excellent student models that outperform teacher models are rare due to inefficient knowledge transfer (Zhang et al.,, 2019).
To address the shortcomings of traditional distillation, we employ the self-distillation framework proposed by Zhang et al., (2019), which adds multiple auxiliary classifiers (AC) to the model. The deepest part and the shallow part with classifiers are respectively treated as the teacher model and the student model, achieving single-stage self-knowledge distillation (SKD). Experimental results show that using SKD improves network performance while maintaining low complexity.
3 Method
The framework of our SKD-STSTAN is illustrated in Fig. 1. Firstly, the onset and apex frames of ME sequences as well as the motionless and real optical flow images are fed into the motion magnification modules to obtain the amplified shape representation, which is discussed in detail in Section 3.1. Then, the magnified apex frame representation is input into the static-spatial stream (S-stream) to obtain the contour and appearance information of the whole face. The representation of the magnified local blocks is input into the local-spatial stream (L-stream) for the local ME feature extraction. The efficient channel attention (ECA) module (Wang et al.,, 2020) used in the L-stream is introduced in Section 3.2. The amplified optical flow maps are input into the dynamic-temporal stream (T-stream), and the temporal shift module (TSM) (Lin et al.,, 2019) used in the T-stream is introduced in Section 3.3. Finally, the outputs of three recognition streams are fused into a fully connected layer to get the classification results. Furthermore, we add multiple auxiliary classifiers (ACs) in the network, leveraging SKD to enable
@@ 3 Method
all blocks to fully learn the features of the training set, which is discussed in Section 3.4.
@@ 3.1 Motion Magnification Module
Due to the low intensity of ME, we enhance muscle movements based on transfer learning using the motion magnification network proposed by Oh et al., (2018). To achieve an end-to-end framework and avoid noise in magnified images reconstructed in the decoder, our motion magnification module does not use a decoder. Given that each stream in our three-stream network has different inputs, we add separate motion magnification modules for each stream. Specifically, the input to the motion magnification network includes two frames with small motion displacement. For an ME sequence, the onset frame is the first frame where the ME begins, and the apex frame is the frame where the facial expression reaches maximum intensity, conveying significant emotional information. Therefore, the onset and apex frames are employed as inputs for the motion magnification module of the S-stream, and the concatenation of their $n\times n$ grid divisions as inputs for the L-stream’s motion magnification module. Particularly, since the T-stream extracts temporal information from optical flow fields, we utilize the optical flow corresponding to no motion and the optical flow between the onset and apex frames (or: between the apex and offset frames) as inputs for the motion magnification module in the T-stream.
Additionally, shape representation often reflects boundaries with geometric properties, while texture representation usually represents color properties (Lei et al.,, 2020). Therefore, shape representation is more helpful for the network to learn the relationships of muscle movements, whereas texture representation provides little useful information for ME. Moreover, the manipulator in the motion magnification module works by calculating the difference between two shape representations and multiplying it by an amplification factor, without using the texture representation, as shown below: